OpenStudy (anonymous):

Solve the following integral using both the general substitution rule and integration by parts:

OpenStudy (anonymous):

\[\int\limits_{0}^{1}e^{-\sqrt(x)}dx\]

OpenStudy (anonymous):

your suggestion disappeared... u= e^(-sqrt(x)) and dv=dx?

OpenStudy (turingtest):

yeah, it wasn't working for me though... maybe it will pay off and I just got lazy

OpenStudy (anonymous):

Yea I haven't gotten very far...

OpenStudy (turingtest):

I'm thinking\[u=e^{-x^{1/2}}\implies du=\frac12x^{-1/2}e^{-x^{1/2}}dx\]\[dv=dx\implies v=x\]\[\int e^{-x^{1/2}}dx=xe^{-x^{1/2}}-\int\frac12x^{1/2}e^{-x^{1/2}}dx\]

OpenStudy (turingtest):

\[u=e^{-x^{1/2}}\implies du=\frac12x^{-1/2}e^{-x^{1/2}}dx\]

OpenStudy (turingtest):

\[\int\frac12x^{1/2}e^{-x^{1/2}}dx=\int xdu=xu-\int udx\]nope, that's the end of it, we get 0=0 from that :P

OpenStudy (turingtest):

new idea: start by substituting \[u=x^{1/2}\]then integrating by parts from the result. I think that will work.